OpenStudy (anonymous):

How would you solve this equation? x+2/5×-10 + x-3/2x-4.

OpenStudy (anonymous):

first combine like terms

OpenStudy (triciaal):

please clarify the original

pooja195 (pooja195):

I think this is what they mean \[\huge~\rm~\bf~\frac{ x+2 }{ 5 \times -10 }+\frac{ x-3 }{ 2x-4 }\]

pooja195 (pooja195):

Im not sure but that's what i think...

OpenStudy (anonymous):

Yes, that's what I meant.